Lucy earned $400 and $550 in interest the last 2 years. Describe the average interest she should earn in each of the next 3 years so that her total earnings over the 5-year period are at least $2300.
Answer by Boreal(15235)

total=2300-400-550=1350 has to be earned in next 3 years
$1350/3=$450 average.